• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Ole Hansen

    Log entry time 16:03:15 on March 28, 2012

    Entry number 367705

    keyword=Day Shift Summary

    08:00 Harp scans in progress
    08:25 No beam, MCC has IOC trouble
    08:40 Beam is back, only to disappear again 5 minutes later for RF rephasing.
    10:15 Beam finally returns. MCC had some unexpected problems before
    rephasing could be done. Continuing harp scans
    11:45 Harp scans done. Moving to target centering study.
    12:50 No beam, injector trouble
    13:15 Beam is back. Resuming target centering study.
    13:35 Target centering done. Moving to optics studies.
    13:40 Left septum trip. Investigating oscillations and spikes seen in
    septa strip charts.
    14:15 Left septum reset. No issues found in either septum aside from the
    mysterious trip. Resuming optics studies.
    14:23 No beam. RF trouble.
    14:44 Beam returns. Resuming optics.
    16:00 Optics studies still in progress